Can you please help me by completing the square?
2x^2-4x-3=0

Answer by stanbon(75887)

Can you please help me by completing the square?
2x^2-4x-3=0
2(x^2-2x+?) =3+2*?
2(x^2-2x+1) = 3 +2*1
2(x-1)^2 = 5
(x-1)^2 = 5/2
x-1 = sqrt(5/2) or x-1 = -(5/2)
x=1+[sqrt10]/2 or x=1-[sqrt10]/2
=========
Cheers,
Stan H.